What is the eccentricity of an ellipse with vertices at (±11, 0) and co-vertices (0, ±5)?
Answer #1

Given Vertices (+1,0) Co-vertices colts) 10,5) (-11,0) (110) ©,-5) Given a=11, b=5 (=a²_b² ² = 112-52=121-25 c² = 96 456 ca r